# Classical Benchmarks for Variational Quantum Eigensolver Simulations of the Hubbard Model

### v1

@@ Abstract
Simulating the Hubbard model is of great interest to a wide range of applications within condensed matter physics, however its solution on classical computers remains challenging in dimensions larger than one. The relative simplicity of this model, embodied by the sparseness of the Hamiltonian matrix, allows for its efficient implementation on quantum computers, and for its approximate solution using variational algorithms such as the variational quantum eigensolver. While these algorithms have been shown to reproduce the qualitative features of the Hubbard model, their quantitative accuracy in terms of producing true ground state energies and other properties, and the dependence of this accuracy on the system size and interaction strength, the choice of variational ansatz, and the degree of spatial inhomogeneity in the model, remains unknown. Here we present a rigorous classical benchmarking study, demonstrating the potential impact of these factors on the accuracy of the variational solution of the Hubbard model on quantum hardware. We find that even when using the most accurate wavefunction ansätze for the Hubbard model, the error in its ground state energy and wavefunction plateaus for larger lattices, while stronger electronic correlations magnify this issue. Concurrently, spatially inhomogeneous parameters and the presence of off-site Coulomb interactions only have a small effect on the accuracy of the computed ground state energies. Our study highlights the capabilities and limitations of current approaches for solving the Hubbard model on quantum hardware, and we discuss potential future avenues of research.
I Introduction
Understanding strongly-correlated electron systems is one of the central challenges of condensed matter physics. Materials with strong electronic interactions exhibit a wealth of interesting properties, including superconductivity [1, 2], Mott insulating behavior [3, 4], excitonic ground states [5, 6], non-Fermi liquid behavior [7], competing and intertwined orders [8, 9, 10], and non-trivial magnetism [11, 12]. One of the most famous models that is capable of capturing such phenomena is the Fermi-Hubbard model [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In its canonical form, its Hamiltonian is written as
@@ I Introduction
$H$ $=\sum_{\sigma}\sum_{\mathbf{R}\mathbf{R}^{\prime}}t_{\mathbf{R}\mathbf{R}^{\prime}}a_{\mathbf{R}}^{\sigma\dagger}a_{\mathbf{R}^{\prime}}^{\sigma}$
@@
Solving for the ground and excited states of the Hubbard model is a quantum many-body problem, which generally scales exponentially with the lattice size when solved exactly on classical computers. As a result, one is restricted to studying finite clusters of relatively small sizes [30], although there exist several approximate methods and numerical techniques which have allowed studying larger systems [31, 32]. On the other hand, the Hamiltonians of Eq. (1) and Eq. (I) are amenable to simulation on quantum computers, for example using variational quantum algorithms (VQAs), where a linear scaling with system size is possible in the ideal scenario [33]. Moreover, several other potential benefits of solving the Hubbard model on quantum computers, as these technologies progress into the fault-tolerant era, have been discussed [34]. These facts make quantum computing an attractive avenue to pursue for tackling such problems. Indeed, there have been several studies solving different flavors of the Hubbard model on quantum computers or classical simulators thereof [35, 36, 37, 38, 39, 40, 41, 42, 43]. Some of these studies have shown that VQAs can reproduce qualitative features of the Hubbard model such as its magnetic properties [39, 40], and it has been shown that for a $1\times 8$ Hubbard chain increasing the ansatz complexity can lead to well-converged values for energies and site occupancies, while spin-spin correlation functions are more challenging to capture [40].
@@
At this point it remains unclear whether the qualitative success of VQAs in describing features of the Hubbard model will also translate into the quantitative reproduction of ground state properties such as the energy, and whether this will depend on the system size and other factors. It is therefore important to systematically benchmark the impact of varying lattice sizes on the quantitative features of the solutions obtained with VQAs, particularly given the importance of extrapolations to the thermodynamic limit for drawing comparisons to established methods [31]. Moreover, it is crucial to establish the impact on VQA simulation accuracy of electronic correlations, the chosen wavefunction ansatz, and the presence of off-site Coulomb interactions and spatially inhomogeneous parameters.
@@
Here we present a detailed classical benchmarking study of the variational quantum eigensolver [44, 45] (VQE), a promising algorithm for quantum simulation on noisy intermediate-scale quantum (NISQ) hardware, applied to single-band Hubbard models of varying degrees of complexity. Our focus lies on comparing the quality of different wavefunction ansätze when scaling to larger system sizes. We perform classical tensor-network simulations of VQE calculations for Hubbard models at half filling on 1D and 2D lattices of up to 16 sites, corresponding to 32 qubits. We consider different onsite interaction strengths $U/t=2,8$ and the inclusion of nearest-neighbor interactions $V$ and spatial inhomogeneities. We show that the so-called number preserving (NP) ansatz, designed specifically for Hubbard models [35], outperforms other popular ansätze, but still incurs errors and converges extremely slowly for larger lattices with strong electronic correlations. Moreover, we demonstrate that VQE with the NP ansatz can successfully describe Hubbard models with nonuniform parameters and interactions beyond the on-site term, which is encouraging for the utility of this approach for solving Hamiltonians representing complex materials. Additionally, we show that an optimization based on maximizing the overlap with a wavefunction computed with a classical reference method can result in a small reduction to the energy error of VQE in some cases, but it can substantially improve the wavefunction fidelity and capture long-range correlations. Our study serves as a reference point along the path to accurate simulation of quantum materials on near-term quantum computers.
II Methods
II.1 Classical simulation of VQE
@@
Our aim is to obtain the ground state of generalized Hubbard models with Hamiltonians of the form in Eq. (I). Here, we focus on models with and without nearest-neighbor repulsion and spatially anisotropic hopping parameters. To do so, we simulate the VQE classically by representing the variational ansatz state as a matrix product state (MPS) within a recently proposed variational tensor network eigensolver (VTNE) approach [46]. Specifically, following Ref. [46], we start from a product state |$\ket{\psi_{0}}$ in a checkerboard configuration, i.e., with alternating spin up and down. We then generate a variational ansatz state via a parameterized quantum circuit (PQC) as follows:
@@
with the Hamiltonian $H$ represented as a matrix product operator (MPO). Within our optimization scheme, we vary the parameters $\bm{\theta}$ of the PQC in order to minimize the energy in Eq. (4). Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}=2N_{x}N_{y}$ is the number of qubits needed to simulate the Hubbard model on an $N_{x}\times N_{y}$ square lattice. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy of the exact PQC. For all tensor operations in this work we have used the ITensor software package [47, 48]. Our VQE results are benchmarked against those obtained within the density matrix renormalization group (DMRG) [49] solution of the studied systems, as implemented within ITensor.
@@
For some cases studied here, we compare the results from this energy-based optimization to those from an overlap-based optimization which minimizes the following loss function, which is defined as the logarithm of the infidelity of the VTNE wavefunction |$\ket{\Psi_{\rm VTNE}}$, with respect to the DMRG wavefunction |$\ket{\Psi_{\rm DMRG}}$, which we consider to be the ground truth:
$f=\log_{10}(1-|\bra{\Psi_{\rm VTNE}}\ket{\Psi_{\rm DMRG}}|^{2}).$ (5)
@@ II.2 Ansätze
II.2.1 Number preserving (NP) ansatz
The number preserving (NP) ansatz [35] is a generalization of the Hamiltonian Variational Ansatz (HVA) [52]. As discussed in Ref. [35], for a 2D square-lattice system, a layer of the HVA applies a unitary operator of the form
@@ II.2.1 Number preserving (NP) ansatz
where $t_{H_{o}},t_{h_{1}},t_{h_{2}},t_{v_{1}},t_{v_{2}}$ are variational parameters. Additionally, $H_{o}$ consists of the on-site terms of the Hamiltonian. The vertical hopping terms in the Hamiltonian are partitioned into components $H_{v_{1}}$,$H_{v_{2}}$ such that each component consists of mutually commuting terms, and the same is done with the horizontal hopping terms, where the partitions are denoted as $H_{h_{1}}$,$H_{h_{2}}$. The NP ansatz replaces each hopping and on-site term within $H_{o}$, $H_{v_{1}}$, $H_{v_{2}}$,$H_{h_{1}}$, and $H_{h_{2}}$, with a general two-qubit NP operator, which takes two parameters $\theta,\phi$:
@@
This generalized HVA introduces independent parameters for each term in the Hamiltonian, unlike the original HVA [52], which introduces a single parameter for each class of terms, e.g., for all horizontal hopping terms. Therefore, while for the NP ansatz used here the number of parameters in the optimization scales with the system size, this is not the case in the original HVA. Since sites that are adjacent in the 2D lattice system are not necessarily adjacent in the 1D qubit encoding, a series of fermionic SWAPs are needed to bring together these sites in the circuit before applying the two-qubit NP operator. After we apply the NP operator, we apply fermionic SWAPs to send the qubits back. As discussed in Ref. [35], a simplified circuit has hopping terms between vertically adjacent qubits implemented locally using an NP gate by applying a parameterless fermionic SWAP gate
@@
A single layer of this ansatz consists of applying the $U_{\text{NP}}$ gate to each pair of qubits coupled by the Hamiltonian of Eq. (1), and requires $10N_{x}N_{y}-4N_{x}-4N_{y}$ parameters for an $N_{x}\times N_{y}$ lattice. Note that the number of parameters in the NP ansatz scales to leading order as $\mathcal{O}(\ell N_{x}N_{y})$, where $\ell$ is the number of HVA layers. In contrast, the number of parameters scales as $\mathcal{O}(\ell)$ in the original HVA of Ref. [33] (see Eq. (6)). Similar to Ref. [46], we apply $R_{z}(\theta)$ gates to each qubit prior to the application of the NP ansatz, which was found to improve optimization.
Figure 1: Energy difference per lattice site, between VQE and DMRG ground state energies when using the NP and EP ansätze, for 1D and 2D lattices, on a logarithmic scale. The shaded region indicates the difference between the minimum and maximum ground state energies obtained using VQE with a total of ten different random starting points. The inset of panel b shows a log-log plot of the minimum number of parameters needed to achieve an error per lattice site $\Delta=0.01$, when using the NP ansatz for 2D lattices.
II.2.2 Excitation preserving (EP) ansatz
We also employ an excitation preserving (EP) ansatz, which is very similar to the NP ansatz outlined above. As implemented within Qiskit [53], we apply the two-qubit gate
@@ II.2.2 Excitation preserving (EP) ansatz
This is very similar to the NP operator of Eq. (7). However, unlike the NP case, we do not perform fermionic swaps between the qubits, but rather apply the gate to neighboring qubits, where for two-dimensional lattices we adopt a one-dimensional indexing scheme, with the index of a site $(x,y)$ given by $(x-1)N_{y}+y$. Therefore, for one-dimensional lattices, the NP and EP ansätze are essentially identical. However, while for two-dimensional lattices the lack of fermionic swaps makes the EP ansatz more straightforward to use than the NP, it also means that vertically adjacent qubits are not coupled directly. As a result, despite using two-qubit gates which are closely related to those of the NP ansatz, the EP ansatz is less tailored to the the physics of the Hubbard model Eq. (1), as will also be reflected by our results on two-dimensional lattices.
II.2.3 Unitary coupled cluster (UCC) ansatz

### v2

@@ Abstract
Simulating the Hubbard model is of great interest to a wide range of applications within condensed matter physics, however its solution on classical computers remains challenging in dimensions larger than one. The relative simplicity of this model, embodied by the sparseness of the Hamiltonian matrix, allows for its efficient implementation on quantum computers, and for its approximate solution using variational algorithms such as the variational quantum eigensolver. While these algorithms have been shown to reproduce the qualitative features of the Hubbard model, their quantitative accuracy in terms of producing true ground state energies and other properties, and the dependence of this accuracy on the system size and interaction strength, the choice of variational ansatz, and the degree of spatial inhomogeneity in the model, remains unknown. Here we present a rigorous classical benchmarking study, demonstrating the potential impact of these factors on the accuracy of the variational solution of the Hubbard model on quantum hardware, for systems with up to $32$ qubits. We find that even when using the most accurate wavefunction ansätze for the Hubbard model, the error in its ground state energy and wavefunction plateaus for larger lattices, while stronger electronic correlations magnify this issue. Concurrently, spatially inhomogeneous parameters and the presence of off-site Coulomb interactions only have a small effect on the accuracy of the computed ground state energies. Our study highlights the capabilities and limitations of current approaches for solving the Hubbard model on quantum hardware, and we discuss potential future avenues of research.
1 Introduction
Understanding strongly-correlated electron systems is one of the central challenges of condensed matter physics. Materials with strong electronic interactions exhibit a wealth of interesting properties, including superconductivity [1, 2], Mott insulating behavior [3, 4], excitonic ground states [5, 6], non-Fermi liquid behavior [7], competing and intertwined orders [8, 9, 10], and non-trivial magnetism [11, 12]. One of the most famous models that is capable of capturing such phenomena is the Fermi-Hubbard model [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In its canonical form, its Hamiltonian is written as
@@ 1 Introduction
$H$ $=-\sum_{\sigma}\sum_{\mathbf{R}\mathbf{R}^{\prime}}t_{\mathbf{R}\mathbf{R}^{\prime}}a_{\mathbf{R}}^{\sigma\dagger}a_{\mathbf{R}^{\prime}}^{\sigma}$
@@
Solving for the ground and excited states of the Hubbard model is a quantum many-body problem, which generally scales exponentially with the lattice size when solved exactly on classical computers. As a result, one is restricted to studying finite clusters of relatively small sizes [32], although there exist several approximate methods and numerical techniques which have allowed studying larger systems [33, 34, 35, 36, 37]. On the other hand, the Hamiltonians of Eq. (1) and Eq. (1) are amenable to simulation on quantum computers, where a linear scaling with system size is possible in the ideal scenario [38]. As quantum computers progress into the (partially) fault-tolerant era, several benefits are expected from solving increasingly complex versions of the Hubbard model [39, 40]. These facts make quantum computing an attractive avenue to pursue for tackling such problems. Indeed, there have been several studies solving different flavors of the Hubbard model on quantum computers or classical simulators thereof [41, 42, 43, 44, 45, 46, 47, 48, 49, 23, 50].
@@
A widely used family of methods for quantum simulation on noisy intermediate-scale quantum (NISQ) hardware [51, 52] are variational quantum algorithms (VQAs) [53], where generally, a trial wavefuction dependent on variational parameters is updated iteratively in order to minimize a cost function. Several studies have shown that VQAs can reproduce qualitative features of the Hubbard model such as its magnetic properties [45, 46], and it has been shown that for a $1\times 8$ Hubbard chain increasing the ansatz complexity can lead to well-converged values for energies and site occupancies, while spin-spin correlation functions are more challenging to capture [46]. However, VQAs are also known to face important challenges, including the barren plateau problem, where minimizing a cost function can become exponentially difficult with the depth of a quantum circuit, due to vanishing gradients [54, 55]. At this point, and given the well-known challenges associated with VQAs, it remains unclear whether the qualitative success of these methods in describing features of the Hubbard model will also translate into the quantitative reproduction of ground state properties such as the energy or correlation functions, and whether this will depend on the system size and other factors. Understanding the quantitative characteristics of the VQA solutions of Hubbard models is particularly timely as the field is transitioning from the NISQ and towards the fault-tolerant era, where systems of increasing complexity may be simulated. It is therefore important to systematically benchmark the impact of varying lattice sizes on the quantitative features of the solutions obtained with VQAs, particularly given the importance of extrapolations to the thermodynamic limit for drawing comparisons to established methods [33]. Moreover, it is crucial to establish the impact of electronic correlations, the chosen wavefunction ansatz, and the presence of off-site Coulomb interactions and spatially inhomogeneous parameters on the accuracy of VQAs, as we are moving towards the quantum simulation of extended Hubbard models that represent real materials.
@@
Here we present a detailed classical benchmarking study of the variational quantum eigensolver [56, 57] (VQE), a promising algorithm for quantum simulation on NISQ hardware [51, 52], applied to single-band Hubbard models of varying degrees of complexity. We perform classical tensor-network simulations of VQE calculations for Hubbard models at half filling on 1D and 2D lattices of up to 16 sites, corresponding to 32 qubits. We consider different ansätze, onsite interaction strengths $U/t=2,8$ and the inclusion of nearest-neighbor interactions $V$ and spatial inhomogeneities. We show that the so-called number preserving (NP) ansatz, designed specifically for Hubbard models [41], outperforms other popular ansätze, but still incurs errors and converges extremely slowly for larger lattices with strong electronic correlations. Moreover, we demonstrate that VQE with the NP ansatz can describe Hubbard models with nonuniform parameters and interactions beyond the on-site term with a similar accuracy to the description of uniform models including only on-site terms, which is encouraging for the utility of this approach for solving Hamiltonians representing complex materials [50]. Additionally, we show that an optimization based on maximizing the overlap with a wavefunction computed with a classical reference method can substantially improve the wavefunction fidelity and capture long-range correlations. Our detailed benchmarking study serves as a reference point along the path to accurate simulation of quantum materials on quantum computers.
2 Methods
2.1 Classical simulation of VQE
@@
Our aim is to obtain the ground state of generalized Hubbard models with Hamiltonians of the form in Eq. (1). Here, we focus on models with and without nearest-neighbor repulsion and spatially anisotropic hopping parameters. To do so, we simulate the VQE classically by representing the variational ansatz state as a matrix product state (MPS) within a recently proposed variational tensor network eigensolver (VTNE) approach [58]. Specifically, following Ref. [58], we start from a product state |$\ket{\psi_{0}}$ in a checkerboard configuration, i.e., with alternating spin up and down. We then generate a variational ansatz state via a parameterized quantum circuit (PQC) as follows:
@@
[…]
Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}=2N_{x}N_{y}$ is the number of qubits needed to simulate the Hubbard model on an $N_{x}\times N_{y}$ square lattice. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy of the exact PQC. For all tensor operations in this work we have used the ITensor software package [59, 60]. Our VQE results are benchmarked against those obtained within the density matrix renormalization group (DMRG) [61] solution of the studied systems, as implemented within ITensor.
@@
For some cases studied here, we compare the results from energy-based optimization to those from an overlap-based optimization, which minimizes the following loss function, defined as the logarithm of the infidelity of the variational wavefunction |$\ket{\psi_{\text{PQC}}(\vec{\theta})}$ with respect to the DMRG wavefunction |$\ket{\psi_{\rm DMRG}}$, considered as the ground truth:
$f=\log_{10}(1-|\bra{\Psi_{\text{PQC}}(\vec{\theta})}\ket{\Psi_{\rm DMRG}}|^{2}).$ (5)
To evaluate the gradient of $f$ with respect to a variational parameter $\theta_{k}$, we again use the intermediate states |$\ket{\psi_{L}^{(k)}}$ and |$\ket{\psi_{R}^{(k)}}$, which isolate the contribution of the gate parameter $\theta_{k}$ during the gradient calculation. The gradient contribution from each gate in the ansatz is computed as:
$\frac{\partial f}{\partial\theta_{k}}=-\frac{2}{IF\ln(10)}$
×$\times\text{Re}\Bigg{(}\bra{\psi_{L}^{(k)}}\partial_{\theta_{k}}U_{k}^{\dagger}\ket{\psi_{R}^{(k)}}\bra{\psi_{\rm DMRG}}\ket{\psi_{\text{PQC}}(\vec{\theta})}\Bigg{)}$ (6)
where $IF=1-|\bra{\Psi_{\text{PQC}}(\vec{\theta})}\ket{\Psi_{\rm DMRG}}|^{2}$. Once again, the intermediate states are updated iteratively, as discussed for the gradient calculation of the energy.
@@ 2.2 Ansätze
2.2.1 Number preserving (NP) ansatz
The number preserving (NP) ansatz is a generalization of the Hamiltonian Variational Ansatz (HVA) [63], and it was designed in Ref. [41] specifically within the context of finding the ground state of the Hubbard model, making it a natural choice for our benchmarking study. As discussed in Ref. [41], for a 2D square-lattice system, a layer of the HVA applies a unitary operator of the form
@@ 2.2.1 Number preserving (NP) ansatz
where $t_{H_{o}},t_{h_{1}},t_{h_{2}},t_{v_{1}},t_{v_{2}}$ are variational parameters. Additionally, $H_{o}$ consists of the on-site terms of the Hamiltonian. The vertical hopping terms in the Hamiltonian are partitioned into components $H_{v_{1}}$,$H_{v_{2}}$ such that each component consists of mutually commuting terms, and the same is done with the horizontal hopping terms, where the partitions are denoted as $H_{h_{1}}$,$H_{h_{2}}$. The NP ansatz replaces each hopping and on-site term with a general two-qubit NP operator, which takes two parameters $\theta,\phi$:
@@
This generalized HVA introduces independent parameters for each term in the Hamiltonian, unlike the original HVA [63], which utilizes a single parameter for each class of terms, e.g., for all horizontal hopping terms. Therefore, while for the NP ansatz used here the number of parameters in the optimization scales with the system size, this is not the case in the original HVA. Specifically, a single layer of this ansatz consists of applying the $U_{\text{NP}}$ gate to each pair of qubits coupled by the Hamiltonian of Eq. (1), and requires $2[2(N_{x}(N_{y}-1)+N_{y}(N_{x}-1))+N_{x}N_{y}]=10N_{x}N_{y}-4N_{x}-4N_{y}$ parameters for an $N_{x}\times N_{y}$ lattice. Note that the number of parameters in the NP ansatz scales to leading order as $\mathcal{O}(\ell N_{x}N_{y})$, where $\ell$ is the number of layers. In contrast, the number of parameters scales as $\mathcal{O}(\ell)$ in the original HVA of Ref. [38] (see Eq. (7)). Similar to Ref. [58], we apply $R_{z}(\theta)$ gates to each qubit prior to the application of the NP ansatz, which was found to improve optimization. This leads to a total of $(10N_{x}N_{y}-4N_{x}-4N_{y})\ell+2N_{x}N_{y}$ variational parameters.
Figure 1: A single layer of the NP ansatz for a $2\times 2$ lattice. Following the application of the on-site terms (purple block), the horizontal hopping terms are implemented (red), followed by fermionic swaps (gray), and subsequently by vertical hopings (blue), and another series of fermionic swaps.
The qubit encoding for the NP ansatz uses a Jordan-Wigner transformation, where each spin-orbital is represented by a qubit. The qubits are indexed using a row-major ordering, where the index of a site $(x,y)$ is given by $(x-1)N_{y}+y$. Since there are two spin species per site, the qubits are ordered with spin-up and spin-down components stored consecutively for each lattice site. This ordering simplifies the circuit construction but introduces a challenge for implementing hopping terms between physically adjacent sites in the 2D lattice, as they may not be directly adjacent in the qubit chain. After we apply the NP operator, we apply fermionic SWAPs to send the qubits back. As discussed in Ref. [41, 64], a simplified circuit has hopping terms between vertically adjacent qubits implemented locally using an NP gate by applying a parameterless fermionic SWAP gate
@@
Even in the 1D case, fermionic SWAP gates are necessary to ensure that hopping only occurs between qubits representing the same spin species, preserving the correct fermionic anticommutation relations. The quantum circuit corresponding to a single layer of the NP ansatz is schmeatically given in Fig. 1, for the case of a $2\times 2$ lattice.
2.2.2 Excitation preserving (EP) ansatz
Figure 2: A single layer of the EP ansatz for a $2\times 2$ lattice.
We also employ an excitation preserving (EP) ansatz, which has certain similarities to the NP ansatz outlined above. As implemented within Qiskit [65], we apply the two-qubit gate
@@ 2.2.2 Excitation preserving (EP) ansatz
This is very similar to the NP operator of Eq. (8). However, unlike the NP case, we simply apply the gate to neighboring qubits according to the indexing scheme where the index of a site $(x,y)$ is given by $(x-1)N_{y}+y$, without each term necessarily reflecting the physical hopping and on-site terms of the Hubbard Hamiltonian. This makes the EP ansatz more generic and straightforward to use compared to NP. However, the EP ansatz is also not tailored specifically to the physics of the Hubbard model on a lattice, as will also be reflected by our results. Within the EP ansatz, no fermionic swaps are performed, as is done within the NP ansatz in an effort to replicate vertical and horizontal hopping terms of the Hubbard Hamiltonian, between same spin qubits. Before and after the application of $U_{EP}$ we apply single-parameter $R_{z}$ gates, leading to a total number of $4N_{x}N_{y}+2(2N_{x}N_{y}-1)\ell$ variational parameters for this generic ansatz. The quantum circuit representing a single layer of the EP ansatz for a $2\times 2$ lattice is visualized in Fig. 2.
2.2.3 Unitary coupled cluster (UCC) ansatz